OpenStudy (anonymous):

M=∫∫∫ρdV Z =∫∫∫(f+φ)ρdV while ρ(x,y,z) is density,M is mass,T is temperature f(1/ρ,T) is helmholtz free energy per unit mass φ(x,y,z) is potential energy per unit mass it's said that when Δz=0 the system is in equilibrium mass doesn't change ΔM=0 then how to solve these two equations so as to make the statement that f+φ+ρ(∂f/∂ρ)=constant , independent of x,y,z ?

OpenStudy (kainui):

How far are you able to get on your own? It looks like we need to use the chain rule in differentiating f with respect to ρ.

OpenStudy (kainui):

Maybe we can also use integration by parts and split Z into two integrals to get it to look like M? Some interesting things to play around with.

OpenStudy (anonymous):

Δ∫∫∫(f+φ)ρdV =∫∫∫[Δ(f+φ)ρ+(f+φ)Δρ]dV =∫∫∫[ρΔf+ρΔφ+(f+φ)Δρ]dV =∫∫∫{ρ[(∂f/∂ρ)Δρ+(∂f/∂T)ΔT]+ρΔφ+(f+φ)Δρ}dV =∫∫∫{[f+φ+ρ(∂f/∂ρ)]Δρ+ρΔφ+ρ(∂f/∂T)ΔT}dV then i don't know how to get any further

OpenStudy (anonymous):

Where, \(\Delta=\nabla^2\)?

OpenStudy (anonymous):

Never mind, that just looks awkward, what do you mean, then, by \(\Delta M=0\)?

OpenStudy (anonymous):

it's calculus of variation.To my knowledge,i think it's about time. the mass doesn't change from time to time i think

OpenStudy (anonymous):

That's what I thought, if you want to make it so, make sure to not use capital delta \(\delta M\) is more appropriate. I'll be right back, see what I can do.

OpenStudy (anonymous):

I must say, I'm currently stumped. It seems like a continuity equation, but I'm not sure what to do with the variations, since it's not a functional we're referring to.

OpenStudy (anonymous):

I mean, we could simply go with the assumption that the system is in equilibrium when, given some small \(|\delta \rho|>0\) we have that \(\delta z=0\). Then it's quite simple.

OpenStudy (anonymous):

well after reading books i find every similar problem it goes that way [f+φ+ρ(∂f/∂ρ)]Δρ then f+φ+ρ(∂f/∂ρ) is constant [g+φ]Δρ then g+φ is constant what i can't figure out is that when it's with Δρ then it's ''coefficient'' must be constant

OpenStudy (anonymous):

does it get something to do with ΔM=∫∫∫ ΔρdV=0 ?

OpenStudy (anonymous):

That's what I'm thinking. Perhaps integration by parts on the first integral; but that gives an odd result.

OpenStudy (anonymous):

What book is this, anyways? It'd be nice to know a bit more about the context.

OpenStudy (anonymous):

well that book is in chinese and it's a bit old

OpenStudy (anonymous):

@dan815

OpenStudy (anonymous):

Again, if we do go by the first definition (which I what I think it means) that for some variation \(\delta \rho \ne 0\), we have that \(\delta z=0\), then the statement is fairly clear. I mean, this is the only thing that makes sense, because any other variation yields unwanted terms, so we'd need stricter conditions to clean them up.

OpenStudy (anonymous):

but what about other terms like ρΔφ+ρ(∂f/∂T)ΔT ?

OpenStudy (anonymous):

All are zero because each is independent of \(\rho\).

OpenStudy (anonymous):

why Δφ is zero? i'm wondering if Δ means the variation from time to time

OpenStudy (anonymous):

In this case, it just means a given variation. \(\delta \phi\) is zero since it is completely independent of density (only dependent on position), and \(\delta T\) is the same.